Question 42
ProbabilityThree boys play a game of luck in which their respective chances of winning are ½, ⅓ and ¼. What is the probability that one and only one of the boys wins the game?
- A.1/24
- B.1/12
- C.11/24
- D.23/24
Show the answer and explanation
Answer: C (11/24)
Explanation
P(exactly one) = ½·⅔·¾ + ½·⅓·¾ + ½·⅔·¼ = 6/24 + 3/24 + 2/24 = 11/24.